# PicoPose: Progressive Pixel-to-Pixel Correspondence Learning for Novel Object Pose Estimation

### v1

@@ III-B Progressive Pixel-to-Pixel Correspondence Learning of PicoPose
With the $L$ offset regression blocks, we have the fine-grained $\mathcal{P}$, with a certainty map $\frac{1}{L}\sum_{i=1}^{L}\mathcal{C}_{l}^{\prime}$. For each foreground pixel in $\mathcal{I}$, if its correspondence certainty exceeds 0.5, we use the position in $\mathcal{P}$ to find the corresponding pixel in $\mathcal{T}$, which is linked to a 3D surface point. Therefore, all the pixel-to-pixel correspondences generate the associated 2D-3D pairs to compute the final object pose via PnP/RANSAC.
III-C Training of PicoPose
@@
where $\mathcal{L}_{coarse}$, $\mathcal{L}_{smooth}$, and $\mathcal{L}_{fine}$ are the loss terms associated with each of the three stages, respectively.
In Stage 1, we adopt the InfoNCE loss [36], as used in GigaPose [20], as the training objective $\mathcal{L}_{coarse}$ to learn feature matching between the RGB observation $\mathcal{I}$ and the best-matched template $\mathcal{T}$.
TABLE I: Quantitative results of different methods on the seven core datasets of BOP benchmark [25]. We report the mean Average Recall (AR) among VSD, MSSD and MSPD, as introduced in Sec. IV-A. ‘FM’ denotes featuremetric pose refinement [21].
Method ##\##Hypothesis BOP Dataset Mean
@@
In Stage 2, we predict the 2D affine transformation $\mathcal{M}$, including in-plane rotation angle $\alpha$, scale $s$, and 2D translation $(t_{u},t_{v})$, to generate smooth correspondences between $\mathcal{I}$ and $\mathcal{T}$. Letting $\hat{\alpha}$, $\hat{s}$, and $(\hat{t}_{u},\hat{t}_{v})$ represent the respective ground truths of the predicted parameters, we define the training objective $\mathcal{L}_{smooth}$ for this stage as follows:
where $\mathcal{L}_{geo}(\alpha,\hat{\alpha})$ is the geodesic distance between two angles $\alpha$ and $\hat{\alpha}$, defined as follows:
$\mathcal{L}_{geo}(\alpha,\hat{\alpha})=\text{acos}\bigl{(}\text{cos}(\alpha)\text{cos}(\hat{\alpha})+\text{sin}(\alpha)\text{sin}(\hat{\alpha})\bigr{)}\>.$ (6)
In Stage 3, we employ the $L_{1}$ distance and the binary cross entropy objective to guide the learning of coordinate offsets $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and certainty maps $\{\mathcal{C}_{l}\}_{l=1}^{L}$ across all $L$ offset regression blocks as follows:
$\mathcal{L}_{fine}=\sum_{l=1}^{L}\lambda||\hat{\mathcal{C}}_{l}\cdot(\Delta\mathcal{P}_{l}-\Delta\hat{\mathcal{P}}_{l})||+\mu\mathcal{L}_{bce}(\mathcal{C}_{l},\hat{\mathcal{C}}_{l}),$ (7)
where $\{\Delta\hat{\mathcal{P}}_{l}\}_{l=1}^{L}$ and $\{\hat{\mathcal{C}}_{l}\}_{l=1}^{L}$ represent the corresponding ground truths of $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and $\{\mathcal{C}_{l}\}_{l=1}^{L}$, while $\lambda$ and $\mu$ are the weights to balance the loss terms. $\mathcal{L}_{bce}$ denotes the binary cross entropy objective. For the supervision of $\Delta\mathcal{P}_{l}$, we use the $\hat{\mathcal{C}}_{l}$ to mask and exclude invalid correspondences from the training process.
IV Experiments
IV-A Experimental Setups
@@
Datasets we train the model on the synthetic datasets of ShapeNet-Objects [23] and Google-Scanned-Objects [24] provided by [17], using a total of $2,000,000$ training images. Evaluation is conducted on the seven core datasets of the BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V.
Implementation Details Following GigaPose [20], we utilize CNOS [22] for zero-shot segmentation and employ $N=162$ templates per object for template matching. The input images are resized to $H\times W=224\times 224$. We use $L=3$ stacked offset regression blocks to refine local correspondence, with feature maps at spatial resolutions of $16\times 16$, $32\times 32$, and $64\times 64$. The hyperparameters $\lambda=0.1$ and $\mu=1.0$ are applied in Eq. (7). PicoPose is trained using the ADAMW optimizer for 400,000 iterations, starting with an initial learning rate of $1\times 10^{-5}$, following a cosine annealing schedule. The training batch size is set to 64, distributed across 8 NVIDIA 3090 GPUs.
@@
Evaluation Metrics We employ the standard BOP evaluation protocol [25] to report the mean Average Recall (AR) w.r.t three error functions, i.e., Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD) and Maximum Symmetry-Aware Projection Distance (MSPD). We also report the End-Point-Error (EPE), a widely used metric in flow estimation [35], to assess the quality of the correspondences.
@@
Figure 4: Qualitative results of different methods without iterative refinement on the seven core datasets of BOP benchmark [25], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V, arranged from left to right.
IV-B Comparisons with Existing Methods
@@
We evaluate our proposed PicoPose against existing methods on the seven core datasets of the BOP benchmark [25]. Inspired by GigaPose [20], we enhance robustness of PicoPose by using the top 5 templates in Stage 2 and Stage 3 to learn fine correspondences and selecting the poses that best match these correspondences. The quantitative results are shown in Table I, where PicoPose significantly outperforms other methods, highlighting its superior zero-shot capability for novel object pose estimation through progressive correspondence learning. For example, a single model of PicoPose using top 5 templates outperforms the single models of GigaPose [20] and FoundPose [21] by $22.4\%$ and $5.1\%$ AR, respectively. In Table I, we also report results with the iterative refinement proposed by MegaPose [17], where PicoPose consistently outperforms the others, whether using 1 or 5 pose hypotheses for refinement. The visualizations in Fig. 4 further validate the advantages of PicoPose.
Results with Single Object Reference Images Since obtaining perfect object CAD models is not always practical, object images are sometimes used as references for pose estimation. Following GigaPose [20], we evaluate PicoPose in the most extreme scenario, i.e., with only one single reference image, by employing Wonder3D [37] to reconstruct the object CAD model from this image for pose estimation. As shown in Table II, PicoPose can successfully handle this extreme setting, achieving results comparable to GigaPose.
Runtime Analysis We report the per-image processing time,including segmentation and pose estimation, of different methods without iterative refinement in Table III. For a fair comparison, both GigaPose [20] and our PicoPose are tested on the same servers; as shown in Table III, PicoPose achieves comparable speeds while delivering more impressive results, demonstrating its accuracy and efficiency. Notably, while FoundPose [21] employs more advanced servers, it still incurs significantly higher computational costs when matching its extensive template library (800 templates per object). In contrast, both PicoPose and GigaPose maintain efficient performance while using only 162 templates per object.
TABLE II: Quantitative results of single reference image on LM-O.
Method ##\##Hypothesis AR
@@
TABLE III: Per-image runtime of segmentation and pose estimation on LM-O. “##\##Hyp” denotes the number of pose hypotheses.
Method ##\##Hyp Server Time (s)
@@
IV-C Ablation Studies and Analyses
We conduct ablation studies on LM-O, T-LESS and YCB-V datasets to evaluate the efficacy of designs in PicoPose. Except for specific cases, the results are achieved using only the best-matched templates in Stage 2 and Stage 3.
TABLE IV: Quantitative comparisons among different stages of correspondence learning.
@@ IV-C Ablation Studies and Analyses
Efficacy of Progressive Correspondence Learning The key to the success of PicoPose lies in its design of progressive pixel-to-pixel correspondence learning. To evaluate this, we first analyze the quality improvements of correspondences and examine their impact on pose estimation. For Stage 2, the predicted affine transformations are applied to obtain correspondences. As shown in Table IV, pose precision improves with finer correspondences, supporting the core claim of this paper. Stage 2 smooths coarse correspondences and filters outliers (Fig. 5), significantly improving translation accuracy (errors $<5$ cm), while its overall 6D pose enhancement over Stage 1 is marginal, as Stage 2 does not include viewpoint rotation updates like Stage 1 with PnP. Stage 3 further enhances precision with a $14.1\%$ AR improvement by locally refining the correspondences from Stage 2.
TABLE V: Quantitative comparisons among different variants of Stage 2.
@@
TABLE VI: Quantitative results with different initial correspondences inputted to Stage 3.
@@
Efficacy of Stage 2 We first assess the effectiveness of Stage 2 by presenting the results without it in Table VI, where we initialize the input correspondences for Stage 3 in two ways: 1) by directly using the coarse correspondences from Stage 1, and 2) by using the predicted poses from Stage 1 via PnP/RANSAC to obtain smoother correspondences. The first approach yields less precise results due to the high noise in the correspondences, while the second is less efficient because of the additional PnP/RANSAC processing. Next, we conduct experimental comparisons with regression based on direct feature concatenation of the observation and the template. As shown in Table V, learning from correspondence maps proves to be more effective, as it explicitly models coarse correspondences and effectively captures the variations in affine transformations. Additionally, we replace our Stage 2 with $\mathbf{F}_{ist}$ from GigaPose [20], which is less efficient, as discussed in Sec. III-B.
Efficacy of Stacked Offset Regression Blocks in Stage 3 In Stage 3, we use $L$ offset regression blocks, specifically $L=3$ in our experiments with feature spatial sizes from DPT [34] set to $16\times 16$, $32\times 32$, and $64\times 64$, to refine correspondences within local regions. The results from different offset regression blocks are reported in Table VII. As shown in the table, performance progressively improves as more blocks are used, indicating that finer correspondences are achieved through progressive local refinement.
Influence of the Number of Templates We follow the setup of GigaPose [20] by using $N=162$ templates per object in our evaluation experiments. In Table VIII, we present additional quantitative results with different numbers of templates for both GigaPose and our proposed PicoPose. The results show improvement as more templates are used, since both methods rely on template matching to select the best-matched template for the target object. However, the rate of improvement slows as the number of templates increases. Notably, PicoPose is more effective than GigaPose when using fewer templates, further highlighting the advantages of PicoPose.
TABLE VII: Quantitative results of different offset regression blocks (denoted as “OR Block”) in Stage 3. “Size” denotes the spatial sizes of aligned feature maps for regression. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
@@
TABLE VIII: Quantitative comparison with GigaPose [20] on the number of templates. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
Method ##\##Temp LM-O T-LESS YCB-V MEAN
@@
V Conclusion
In this paper, we propose PicoPose, a novel framework for object pose estimation from RGB images that uses progressive pixel-to-pixel correspondence learning across three carefully designed stages. We demonstrate the zero-shot capabilities of PicoPose on seven core datasets of the BOP benchmark. In future work, we aim to further increase the speed of PicoPose to achieve real-time performance and explore ways to reduce its reliance on templates.

### v2

@@ 3.2 Progressive Pixel-to-Pixel Correspondence Learning of PicoPose
With $L$ offset regression blocks, we have the fine-grained $\mathcal{P}$, with a certainty map $\frac{1}{L}\sum_{i=1}^{L}\mathcal{C}_{l}^{\prime}$. For each foreground pixel in $\mathcal{I}$, if its correspondence certainty exceeds 0.5, we use the position in $\mathcal{P}$ to find the corresponding pixel in $\mathcal{T}$, which is linked to a 3D surface point. Therefore, all the pixel-to-pixel correspondences generate the associated 2D-3D pairs to compute the final object pose.
3.3 Training of PicoPose
@@
where $\mathcal{L}_{coarse}$, $\mathcal{L}_{smooth}$, and $\mathcal{L}_{fine}$ are the loss terms associated with each of the three stages.
In Stage 1, we adopt the InfoNCE loss [36], as used in GigaPose [21], as the training objective $\mathcal{L}_{coarse}$ to learn feature matching. In Stage 2, we predict the 2D affine transformation $\mathcal{M}$, including in-plane rotation angle $\alpha$, scale $s$, and 2D translation $(t_{u},t_{v})$, to generate smooth correspondences between $\mathcal{I}$ and $\mathcal{T}$. Letting $\hat{\alpha}$, $\hat{s}$, and $(\hat{t}_{u},\hat{t}_{v})$ represent the respective ground truths of the predicted parameters, we define the training objective $\mathcal{L}_{smooth}$ for this stage as follows:
[…]
In Stage 3, we employ the $L_{1}$ distance and the binary cross entropy objective to guide the learning of coordinate offsets $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and certainty maps $\{\mathcal{C}_{l}\}_{l=1}^{L}$ across all $L$ offset regression blocks:
[…]
where $\{\Delta\hat{\mathcal{P}}_{l}\}_{l=1}^{L}$ and $\{\hat{\mathcal{C}}_{l}\}_{l=1}^{L}$ represent the corresponding ground truths of $\{\Delta\mathcal{P}_{l}\}_{l=1}^{L}$ and $\{\mathcal{C}_{l}\}_{l=1}^{L}$, while $\lambda$ and $\mu$ are the weights to balance the loss terms. $\mathcal{L}_{bce}$ denotes the binary cross entropy objective. For the supervision of $\Delta\mathcal{P}_{l}$, we use $\hat{\mathcal{C}}_{l}$ to mask and exclude invalid correspondences.
4 Experiments
Method #\#Hypothesis BOP Dataset Mean
@@ 4 Experiments
Table 1: Quantitative results of different methods on BOP datasets [26]. We report the mean Average Recall (AR) among VSD, MSSD and MSPD. ‘FM’ denotes featuremetric pose refinement [22].
@@
Datasets. we train PicoPose on synthetic datasets of ShapeNet-Objects [24] and Google-Scanned-Objects [25] provided by [18], using a total of 2 million training images. Evaluation is conducted on seven BOP datasets [26], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V.
@@
Evaluation Metrics. We report the mean Average Recall (AR) w.r.t three error functions, i.e., Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) [26] . We also report the end-point-error (EPE), a widely used metric in flow estimation [35], to assess the quality of the correspondences.
@@
Figure 4: Qualitative results of different methods without iterative refinement on BOP datasets [26], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V, arranged from left to right.
4.1 Comparisons with Existing Methods
@@
We evaluate PicoPose against existing methods on the seven core datasets of the BOP benchmark [26]. Inspired by GigaPose [21], we enhance the robustness of PicoPose by using the top 5 templates in Stage 2 and Stage 3 to learn fine correspondences and selecting the poses that best match these correspondences. The quantitative results are shown in Table 1, where PicoPose significantly outperforms other methods, highlighting its superior zero-shot capability for novel object pose estimation through progressive correspondence learning. For example, a single model of PicoPose using top 5 templates outperforms the single models of GigaPose [21] and FoundPose [22] by $22.4\%$ and $5.1\%$ AR, respectively. In Table 1, we also report results with the iterative refinement proposed by MegaPose [18], where PicoPose consistently outperforms the others, whether using 1 or 5 pose hypotheses for refinement. The visualizations in Fig. 4 further validate the advantages of PicoPose.
Method #\#Hypothesis AR
@@
Table 2: Quantitative results of single reference image on LM-O.
Method #\#Hyp Server Time (s)
@@
Table 3: Per-image runtime on LM-O.
Results with Single Object Reference Images. Since obtaining perfect object CAD models is not always practical, object images are sometimes used as references. Following GigaPose [21], we evaluate PicoPose in the most extreme scenario, i.e., with only one reference image, by employing Wonder3D [37] to reconstruct the object CAD model from this image. As shown in Table 3, PicoPose can successfully handle this extreme setting, achieving results comparable to GigaPose.
Runtime Analysis. We report the per-image processing time, including segmentation and pose estimation, of different methods without iterative refinement in Table 3. For a fair comparison, both GigaPose [21] and PicoPose are tested on the same servers; as shown in Table 3, PicoPose achieves comparable speeds while delivering more impressive results, demonstrating its accuracy and efficiency. Notably, while FoundPose [22] employs more advanced servers, it still incurs significantly higher computational costs when matching its extensive template library (800 templates per object). In contrast, both PicoPose and GigaPose use only 162 templates per object.
@@
Table 4: Quantitative comparisons among different stages of correspondence learning.
4.2 Ablation Studies and Analyses
We conduct ablation studies to evaluate the efficacy of designs in PicoPose. Except for specific cases, the results are achieved using only the best-matched templates in Stage 2 and Stage 3.
@@ 4.2 Ablation Studies and Analyses
Table 5: Quantitative comparisons among different variants of Stage 2.
Efficacy of Progressive Correspondence Learning. The key to the success of PicoPose lies in its design of progressive pixel-to-pixel correspondence learning. To evaluate this, we first analyze the quality improvements of correspondences and examine their impact on pose estimation. For Stage 2, the predicted affine transformations are applied to obtain correspondences. As shown in Table 4, pose precision improves with finer correspondences, supporting the core claim of this paper. Stage 2 smooths coarse correspondences and filters outliers (Fig. 5), significantly improving translation accuracy (errors <5<5 cm), while its overall 6D pose enhancement over Stage 1 is marginal, as Stage 2 does not include viewpoint rotation updates like Stage 1 with PnP. Stage 3 further enhances precision with a $14.1\%$ AR improvement by locally refining the correspondences from Stage 2.
Efficacy of Stage 2. We first assess the effectiveness of Stage 2 by presenting the results without it in Table 6, where we initialize the input correspondences for Stage 3 in two ways: 1) by directly using the coarse correspondences from Stage 1, and 2) by using the predicted poses from Stage 1 via PnP/RANSAC to obtain smoother correspondences. The first approach yields less precise results due to the high noise in the correspondences, while the second is less efficient because of the additional PnP/RANSAC processing. Next, we conduct experimental comparisons with regression based on direct feature concatenation of the observation and the template. As shown in Table 5, learning from correspondence maps proves to be more effective, as it explicitly models coarse correspondences and effectively captures the variations in affine transformations. Additionally, we replace our Stage 2 with $\mathbf{F}_{ist}$ from GigaPose [21], which is less efficient, as discussed in Sec. 3.2.
Efficacy of Stacked Offset Regression Blocks in Stage 3. In Stage 3, we use $L$ offset regression blocks, specifically $L=3$ in our experiments with feature spatial sizes from DPT [34] set to $16\times 16$, $32\times 32$, and $64\times 64$, to refine correspondences within local regions. The results from different offset regression blocks are reported in Table 7, where performance progressively improves as more blocks are used, indicating that finer correspondences are achieved through progressive local refinement.
@@
Table 6: Quantitative results with different initial correspondences inputted to Stage 3.
@@
Table 7: Quantitative results of different offset regression blocks (denoted as “OR Block”) in Stage 3.
5 Conclusion
In this paper, we propose PicoPose, a novel framework for object pose estimation from RGB images that uses progressive pixel-to-pixel correspondence learning across three carefully designed stages. We demonstrate the zero-shot capabilities of PicoPose on seven core datasets of the BOP benchmark, enabling practical deployment in robotic applications. In future work, we aim to further increase the speed of PicoPose to achieve real-time performance and explore ways to reduce its reliance on templates.
6 Limitations
While PicoPose demonstrates strong performance on 6D object pose estimation benchmarks with real-world cluttered scenes and shows promising potential for rapid deployment in robotic applications through simulated grasping experiments, several areas remain for future improvement.
First, the reliance on multiple templates for 3D object representation means that its performance is inherently tied to the number of templates used. Although our approach achieves comparable results with fewer templates than existing methods (as detailed in Supplementary Section 2), future work could explore more efficient template utilization strategies or alternative 3D representations.
Second, while PicoPose benefits from the iterative refinement of MegaPose [18], the performance gains are relatively modest compared to other methods, probably because PicoPose’s initial estimates already approach the refiner’s performance upper bound. This motivates the development of specialized refinement approaches for high-quality initial predictions.
Third, in case of a single reference image of the target object, where we follow GigaPose [21] to use Wonder3D for reconstructing the object in 3D space, the quality of reconstruction remains a limiting factor that affects final pose estimation accuracy. We believe integrating emerging neural reconstruction techniques could significantly improve performance in this challenging scenario.
Appendix
This appendix is structured as follows. In Section A, we provide more details on the network architecture, training setup, equipment, pose solving configuration, and data augmentation methods used in this work. In Section B, we perform more ablation studies on the effects of correlation lookup in Stage 3 and the influence of the number of templates on pose precision. Section C shows more qualitative results of different stages in PicoPose and more qualitative comparisons of different methods on real-world benchmarks. In Section D, we conduct robotic grasping experiments in a simulation environment using the PyBullet physics engine [38] to verify the application of our method in robotic grasping.
Appendix A More Experimental Details
Network architecture details. In Stage 1, we build on previous work [21] by using dinov2_\_vitl14 as our backbone. The feature dimension of this backbone is $1,024$ for each token. In Stage 2, we employ two conventional layers with group normalization and a ReLU activation function to reduce the spatial size of the input correspondence map $\mathcal{A}$ to $8\times 8$, after which we flatten the feature map to obtain the global pose vector. In Stage 3, in order to reduce network parameters, we set the number of channels $D_{l}$ of the $l^{th}$ feature map generated from DPT [34] to 256. Additionally, we utilize the standard lookup operation in RAFT [35] for $L$ blocks. Since the feature sizes in each block vary, the hyperparameter settings of each block need to be adjusted individually. For the $l^{th}$ block, we establish the layer of the correlation pyramid as $l+1$ and set the radius of the correlation lookup to 4. We also list the model sizes of different stages in Table 8.
Stage #Param
1 304 M
2 18 M
3 58 M
Total 380 M
Table 8: The model sizes of different stages in our network.
Details of training settings. We report the detailed hyperparameter settings to train our network in Table 9.
Hyperparameters Settings
Optimizer AdamW
AdamW $\beta$ (0.5, 0.99)
AdamW epseps 1e-6
Learning rate scheduler Cosine decay
Training iterations $400,000$
Warmup iterations $1,000$
Learning rate 1e-5
Weight decay 5e-4
Batch size 32
Table 9: Detailed hyperparameters in training our network.
More details of devices. We conducted all experiments with GPU in GeForce RTX 3090 24G, and CPU in Intel (R) Xeon (R) CPU E5-2678 v3 @ 2.50 GHz under the Linux operating system.
Details of PnP/RANSAC. We utilize the EPnP algorithm [39] along with the RANSAC scheme in the fine correspondence to solve the object pose in Stage 3. The RANSAC iterations are configured to 150, and the reprojection error threshold is set to 2.
Details of augmentations. In Stage 1, to better adapt to the input images in real-world scenarios, we conduct data augmentation on the query image of the training data in a manner similar to GDRNPP [40]. In Stage 3, we apply random 2D translation, in-plane rotation, and scale noise to the ground truth affine transformation $\mathcal{M}$ to generate the initial coordinate map $\mathcal{P}$.
Appendix B More Ablation Studies
Effects of correlation lookup in Stage 3. In Stage 3, we use the correlation lookup operation in RAFT [35] to obtain flow features, and combine them with the features of the input image and the best-matched template to predict coordinate offsets and the certainty map. To verify the effectiveness, we conducted experiments without using the correlation lookup operation. As shown in Table 10, the features obtained by the correlation lookup operation can improve the results significantly.
Correlation Lookup LM-O T-LESS YCB-V MEAN
$\times$ 35.0 25.9 46.8 35.9
46.3 39.7 58.7 48.2
@@ Appendix B More Ablation Studies
Table 10: Quantitative results of correlation lookup operation in Stage 3. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
Influence of the number of templates. We follow the setup of GigaPose [21] by using $K=162$ templates per object in our evaluation experiments. In Table 11, we present additional quantitative results with different numbers of templates for both GigaPose and our proposed PicoPose. The results show improvement as more templates are used, since both methods rely on template matching to select the best-matched template for the target object. However, the rate of improvement slows as the number of templates increases. Notably, PicoPose is more effective than GigaPose when using fewer templates, further highlighting the advantages of PicoPose.
Method #\#Temp LM-O T-LESS YCB-V MEAN
@@
Table 11: Quantitative comparison with GigaPose [21] on the number of templates. We report the mean Average Recall (AR) among VSD, MSSD and MSPD.
Appendix C Additional Qualitative Results
More qualitative results of different methods. We present more qualitative results of different methods on the seven core datasets (LM-O[41], T-LESS[42], TUD-L[43], IC-BIN[44], ITODD[45], HB[46], and YCB-V[6] ) in the BOP benchmark [26], shown in Fig. 6. We illustrate the estimated 6D pose by rendering the 3D model on the input image and using the overlap rate as a basis, where a higher overlap rate indicates a more accurate estimated 6D pose. Specifically, all methods use the same zero-shot segmentation method, i.e., CNOS [23].
Figure 6: More qualitative results of different methods on the seven core datasets of BOP benchmark [26], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V, arranged from top to bottom.
More qualitative results of different stages. We visualize the correspondences between the query image $\mathcal{I}$ and the best-matched template $\mathcal{T}$ in Stage 1 and Stage 2. As shown in Fig. 7, the coarse correspondences generated in Stage 1 contain numerous outliers and inconsistencies, many of which are effectively resolved by Stage 2 to produce smooth correspondences. In Stage 3, we enhance the display of fine correspondences by visualizing the coordinate map $\mathcal{P}$ as optical flow with the certainty map on YCB-V dataset [6], shown in Fig. 8.
Figure 7: Qualitative results of coarse correspondences in Stage 1 and smooth correspondences in Stage 2 on the seven core datasets of BOP benchmark [26], including LM-O, T-LESS, TUD-L, IC-BIN, ITODD, HB, and YCB-V, arranged from top to bottom.
Figure 8: Qualitative results of fine correspondences in Stage 3 on YCB-V dataset [6].
Appendix D Application: Robotic Grasping in Simulated Environments
In this section, our proposed PicoPose demonstrates seamless integration for robotic grasping applications using PyBullet [38] with the setup shown in Fig. 9. Our experimental scene comprises (1) a Franka Emika Panda robotic arm, (2) distractor objects, including the target, randomly arranged on the workspace, and (3) a placement tray. A fixed virtual camera captures single RGB images of the cluttered scene as input to our system.
The processing pipeline consists of three key stages. First, CNOS [23] segments the target object in the RGB scene. Second, our proposed PicoPose estimates the 6D pose of the target object in camera coordinates. Third, we use the known camera-to-robot coordinate transformation to convert this pose into a 6D grasping pose and use inverse kinematics to generate robot motions to successfully grasp the target object.
Figure 9: Robotic grasping application of PicoPose in simulated environment.